prove that if the sum of the sequence of the roots of the equation ax^2+bx+c=0 is 1 then b^2=2ac=a^2
Answer by ikleyn(52788)

.
prove that if the sum of the sequence of the roots of the equation ax^2+bx+c=0 is 1 then b^2=2ac=a^2
~~~~~~~~~~~~~~~~


            Let's check if the statement of the post is valid.


Let's consider the equation


    x*(x-1) = 0,  or,  equivalently,  x^2 - x = 0.


It has the roots  x= 0  and  x= 1;  the sum of the roots is 1,  which is in accordance with the description.


The coefficients of the standard form equation are  a = 1, b = -1, c = 0.


The statement asserts that  a^2 = b^2, and it is true.


It asserts also that a^2 = 2ac, but it is  WRONG :  a^2 = 1, 2ac = 0  and 1 =/= 0.


Thus the problem' statement is  FALSE.


In other words, this "quasi"-problem is a  FAKE.

----------

It seems to me,  that I see this problem not for the first time at this forum.

Let me  AWARE  you,  that  -EITHER-  the source of the problem is  UNTRUSTED,  -OR-  you incorrectly reproduce it from the source.


            - - - Or     B O T H - - - .


In any case,  by posting it,  you  STOLE  the tutors'  valuable time.


If I see it next time at the forum,  I will delete it without any explanations.



===========

At this spot,  a real problem from a real student could be placed,  who really needs my help and my attention.

And during this time,  that I spent on your post  FOR  NOTHING,  I could help at least to two real students.